OpenStudy (anonymous):

Solve the equation by factoring. 12x^2+12x=24 x= (smaller value) x= (larger value)

OpenStudy (anonymous):

12x^2+12x=24 12x^2+12x-24=0 12(x^2+x-2)=0 12(x-1)(x+2)=0 x=1 larger and x=-2 smaller value